### v1

@@ Abstract
We extend the construction of Alcubierre–Natário class of warp drives to an infinite class of spacetimes with similar properties. This is achieved by utilising the Martel–Poisson charts which closely resembles the Weak Painlevé–Gullstrand form for various background metrics (Mink, AdS, dS). The highlight of this construction is the non–flat intrinsic metric which in three dimensional spacetimes introduces conical singularities at the origin and in higher dimensions generates a non–zero Ricci scalar for the spatial hypersurfaces away from the origin. We analyse the expansion/contraction of space and $(-ve)$ energy densities associated with this class of warp drives and find interesting deviations due to the global imprints of the conical defects. Several generalizations are also discussed.
@@
Warp Drive Conical defects Martel–Poisson ChartsPainlevé–Gullstrand
@@ 1 Introduction
In this paper we have generalized the construction of Alcubierre like warp drive in three dim Minkowski spacetime by leveraging the Martel–Poisson (MP) family of charts Martel_2001 of which the Painlevé–Gullstrand (PG) coordinates are a special case. This construction has been extended to $\text{AdS}_{3}$, dS backgrounds and their conical defect counterparts. A generic key feature is the appearance of a non–flat intrinsic metric with conical singularity which affect the global properties of the spacetime. Another important realization comes from the flow velocities of the background metrics where the expansion and $(-ve)$ energy density of the warp drive is considerably reduced if the drive travels along the background flow velocities. This will be useful in analog gravity setups which imitate the warp metric Fischer_2003 ; finazzi2012analogue and can be studied experimentally PhysRevLett.105.240401 . Throughout the paper we have used natural units $G=c=1$ and metric signature (,+,+,)(-,+,+$\cdots$)( - , + , + , ⋯ ).
The rest of the paper is organized as follows. In section 2 we review and extend the construction of MP–charts for spacetimes of the form (1) in $D$ dimensions. Interpretation of the flow coordinates are discussed including the special case of PG–coordinates. Section 3 highlights the conical singularities of spatial hypersurfaces in three dimensional Minkowski and AdS spacetimes and their conical defect cousins. Regularization of the singularities and embedding of the warp drive in these spacetimes are discussed. We conclude in section 4 with comments on possible explorations to higher dimensions, analogue gravity, quantum effects and optimizations. The paper ends with the relevant references.
@@ 2 Martel–Poisson charts
where $d\Omega_{(D-2)}^{2}$ is the line element on an unit $(D-2)$ sphere and $f(r)=(1+\tilde{f}(r))$  [^1] See later sections 3 and 4 for generalizations.. The main idea is to describe the spacetime in terms of radially infalling or outgoing timelike geodesic observers with D–velocities $u^{\mu}=\frac{dx^{\mu}}{d\tau}=(\dot{t},\dot{r},0,\cdots,0)$ where $\tau$ is the proper time along the geodesic. For static spacetimes, the Killing vector $\xi=(\partial/\partial t)$ leads to a conserved energy per unit mass $\epsilon=-\xi_{\mu}u^{\mu}=f\dot{t}$ and utilizing the normalization of the D–velocities $u_{\mu}u^{\mu}=-f\dot{t}^{2}+(\dot{r}^{2}/f)=-1$, we can finally write
@@
where the ($-ve$) and ($+ve$) signs are for the infalling and outgoing geodesics respectively. So far, we have not established that these D–velocities are indeed geodesics. We demand that this congruence of D–velocities are hypersurface orthogonal to timelike surfaces $T(t,r)$ = constant i.e. $u_{\mu}=-(\partial T/\partial x^{\mu})$ where $T$ is now the proper time for these free falling (geodesic) observers. Redefining the time coordinate as  [^2] Note that $T$ is an exact differential i.e. $\frac{\partial\epsilon}{\partial r}=\frac{\partial\left(\frac{\sqrt{\epsilon^{2}-f}}{f}\right)}{\partial t}\,$.
@@
Note that except for the case $\epsilon=1$, the induced metric on the hypersurface is not Ricci flat. In fact, as we will see later the scalar curvature and other curvature invariants of the induced metric blows up at $r=0$.
@@
The observers can start at infinity but now with some initial velocity $v_{\infty}$ w.r.t. time coordinate $t$ such that $\epsilon=\frac{1}{\sqrt{1-v_{\infty}^{2}}}\geq 1$. Again, zero initial velocity gives the PG–coordinates for the Schwarzschild black holes.
There is another interesting limit where we can set $v_{\infty}=1$ i.e. the speed of light. Scaling $\tilde{T}=(T/\epsilon)$ in (3) and then taking the limit $\epsilon\rightarrow\infty$ gives the infalling Eddington–Finkelstein (EF) coordinates Faraoni2020WhenPC .
Of course, it is also possible to have finite radius non–zero initial velocity observers but for the remainder of this paper we will focus on the initial velocity interpretation of the parameter $\epsilon$.
@@ 3 Warp Drive in various spacetimes
While the original Alcubierre warp drive is a deformation of the Minkowski spacetime, we will now discuss embedding the drive in various spherically symmetric static spacetimes, see also garattini2023black ; garattini2024wormholewarp . Our focus will be on the minimal non–trivial dimension $D=3$ from which it is relatively clear how to generalize for higher dimensions. Three dimensions is also special as it admits Conical defect solutions which change the global properties of the spacetime retaining the local ones.
@@ 3.1 Minkowski spacetime
The MP–charts for the Minkowski spacetime in three dimensions with a tunable parameter $\epsilon$ takes the form,
@@
where the radially infalling geodesic observers have the velocity $v_{\infty}=\sqrt{1-\frac{1}{\epsilon^{2}}}$ at spatial infinity. Though this metric has non–vanishing Christoffel symbols, it is just an artifact of the coordinate choice and the curvature tensors vanish.
@@
It is instructive to analyze the induced metric on the $T$ = constant hypersurfaces,
@@
We can however, introduce a regularization by rolling off the cone tip. The simplest one corresponds to changing $\mathcal{C}_{\epsilon}$ to a hyperbolic space with a tunable regularization parameter $a$,
@@
To evaluate the curvature scalar, we first evaluate the integral curvature on the regularized hyperbolic space in the $a\rightarrow 0$ limit,
@@
where the second integral is zero. In a more general setting one can embed the conical defect in a non Ricci flat two dimensional manifold and the second integral would collect the curvature around the defect. We note that as only the singular point $\tilde{r}=0$ can give rise to the first term in (12), it is best to introduce a local representation for the curvature on $C_{\epsilon}\,$,
@@
We can embed the Alcubierre warp drive in the infalling Minkowski spacetime background (7). It is crucial to note that this embedding can’t be achieved by a coordinate transformation of the warp drive in the usual $\epsilon=1$ charts. Therefore, it represents an infinite class of different embeddings parameterized by $\epsilon$ or equivalently $v_{\infty}$.
@@
where $N$ is the lapse, $N^{i}$ are the shift vectors and $\gamma_{ij}$ are the components of the induced metric on the $T$ = constant timelike hypersurfaces. For the metric as in (7), the embedding is as follows,
@@
where for three dimensions, $r=\sqrt{x^{2}+y^{2}}$ is the radial coordinate, $r_{s}$ is the distance from the center of the drive which is moving along the ($+ve$) $x$–direction with constant velocity $v$. Here, the form function $f(r_{s})$ decides the size and shape of the wrap drive bubble, its exact form is of little consequence (unless one is optimizing) as long as it has the value $f(0)=1$ at the center of the bubble, $f\rightarrow 0$ as $r\rightarrow\infty$ outside the bubble and drops sharply from $1$ to 00 at the wall of the bubble at some distance $R$ from the center of the bubble  [^3] One such specific function popular in the literature is $f(r_{s})=\frac{\tanh[\sigma(r_{s}+R)]-\tanh[\sigma(r_{s}-R)]}{2\tanh(\sigma R)}\,,$ (19) where $\sigma$ has the interpretation of being inversely proportional to the bubble wall thickness and can be taken to be large.. The choice of shift vectors $N^{i}$ are inspired by ellis2004warp , where the flow velocity is that of the warp drive inside the bubble and that of the background spacetime outside the bubble. The Eulerian observers with geodesics $u_{\mu}=(-1,0,0)$ are still in free fall and hence the above metric can be considered as a warp drive different from the Natário type Nat_rio_2002 . We assume the warp drive to be a test particle moving along the timelike curve $x=x_{0}(T)=vT$, regardless of the value of $v(T)$. One can verify that the proper time along this curve equals the coordinate time. Note that here, the induced metric is not flat and as discussed earlier will exhibit conical singularity at the origin. Warp drives with non flat $\gamma_{ij}$ have been discussed earlier in Broeck_1999 ; garattini2024wormholewarp . As expected, for $\epsilon=1$, we get back the original Alcubierre warp drive Alcubierre_1994 .
@@
Few comments on the distance $r_{s}$ and the induced metric $\gamma_{ij}$ are in order. We can always choose $r_{s}$ to be the coordinate distance $r_{s}=\sqrt{(x-x_{0}(T))^{2}+y^{2}}$ from the center of the warp drive. In general, the induced metric $\gamma_{ij}$ is not $\delta_{ij}$, therefore the size & shape of the bubble inside the boundary $r_{s}=R$ will evolve over time. Even without the warp drive, the metric in (7) has non–zero expansion/contraction and shear for a spacelike patch orthogonal to the congruence of Eulerian geodesic observers  [^4] Non zero rotation is not possible as Eulerian geodesics are hypersurface orthogonal.. Our focus will be on the large $x_{0}$ and $(\epsilon-1)$ small limit where these background expansion and shear are at check. However, there are other possible choices, either $r_{s}=\sqrt{\gamma_{ij}(x-x_{0}(T))^{i}(x-x_{0}(T))^{j}}$ or $r_{s}=\text{Geodesic}(x,x_{0}(T))$ where $\text{Geodesic}(x,x_{0}(T))$ is the geodesic distance between the center of the warp drive $(x_{0}(T),0)$ and some point $(x,y)$. Another possible choice for the induced metric is $\gamma_{ij}=\delta_{ij}+(1-f(r_{s}))\left(\frac{1}{\epsilon^{2}}-1\right)\frac{x^{i}x^{j}}{r^{2}}\,$, an interpolating metric which is $\delta_{ij}$ inside the bubble and $\gamma_{ij}$ as in (17) outside  [^5] It is difficult to implement the other choices of $r_{s}$ with this choice of induced metric as now the definition of $r_{s}$ depends on $r_{s}$ itself through the function $f(r_{s})$. This puts severe constraints on possible $f(r_{s})$, which we checked to be incompatible with the desired properties of $f(r_{s})$.. We will discuss some of these cases in A.
@@
With our first choice of $r_{s}$ and the induced metric (17) we shall now discuss the movement of the warp bubble due to the expansion and contraction of the space around it. The space expands behind the warp bubble and contracts in front of it, thereby pushing the bubble forward (and maybe a spaceship inside it) at arbitrary velocity $v$  [^6] The same mechanism also applies to the the expanding universe, wherein galaxies move faster than light w.r.t. each other due to the expansion of the space itself.. We shall also discuss the energy density as measured by the Eulerian observers i.e. $\rho=T_{\mu\nu}u^{\mu}u^{\nu}$ where $T^{\mu\nu}$ is the stress energy tensor of the metric and $u^{\mu}=(1,-N^{x},-N^{y})$ are the 3–velocities.
@@
where $D_{i}$ are the covariant derivatives w.r.t. the induced metric $\gamma_{ij}$. The expansion is the trace $\theta=K^{i}_{i}$. The Einstein tensor $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}$ along with the Gauss–Codazzi relations which connects curvature tensors of the metric to curvatures on the hypersurface(s) give,
@@
$D_{j}K^{j}_{i}-D_{i}K=8\pi p_{i}\,,$
where $\rho$ is the matter energy density and $p_{i}$ are the matter momentum density and $\Lambda$ is the cosmological constant.
For the background metric in (7) without the warp drive, the extrinsic curvature and expansion away from the conical singularity at $r=0$ are given by  [^7] The regularization mentioned in (11) needs modification as now we want ${}^{(2)}R+\left(K^{2}-K^{ij}K_{ij}\right)$ to vanish at $r=0$ as $\rho=0$. The regularized metric around $r=0$ with $0\leq\phi<2\pi$ is given by $ds^{2}=-dT^{2}+\frac{1}{\epsilon^{2}}\left(A(r)\,dr+\sqrt{\epsilon^{2}-1}\,B(r)\,dT\right)^{2}+r^{2}d\phi^{2}$ $A(r)=\frac{a^{2}\epsilon^{4}+r^{2}}{a^{2}\epsilon^{2}+r^{2}}\quad B(r)=\frac{\sqrt{\epsilon^{2}(1-\log(A(r)))-1}}{\sqrt{\epsilon^{2}-1}}$ (22)   .
@@
Closed form expressions for the $K_{ij}$ and $\rho$ can be obtained for the metric with the warp drive centered at $x_{0}$. However, we are interested in the limit $x_{0}\gg(x-x_{0})\approx y\approx r_{s}\,$, where the warp drive is sufficiently far way from the center $r=0$. We shall also take $v_{\infty}\ll 1$ for the Eulerian observers. To $\mathcal{O}(\frac{1}{r})$ and $\mathcal{O}(v_{\infty})$, the expressions are as follows  [^8] Here, ${}^{(2)}R=0$ on the hypersurfaces, except at $r=0$. This is not true in $D\geq 4$ dimensions, thereby opening up new possibilities to minimize the energy density.
@@
At first, we note that for $v_{\infty}=0$, the expressions falls back to the usual Alcubierre warp drive in Minkowski spacetime. As expected, the zeroth order expressions for both $\theta$ and $\rho$ have the combination $(v+v_{\infty})$ for velocity. In these river flow coordinates, the background metric is flowing like a fluid along the radially inward direction. However, the wrap drive is flowing against this “fluid” in the $(+ve)$ $x$–direction, i.e. it is flowing with velocity $(v+v_{\infty})$ w.r.t. the fluid. Therefore, it is best to move along the flow to minimize the peak expansions/contractions and $(-ve)$ energy density. In fact, for a free falling warp drive to zeroth order they are zero. We also note that, the first $\mathcal{O}\left(\frac{1}{r}\right)$ term in $\theta$ is that of the background metric and maybe subtracted, the other one is proportional to $(y/r)$ and hence very small. By construction $f^{\prime}(r_{s})$ is non zero only along a very small width of the ring $r_{s}=R$. Hence, the effects of the warp drive in both $\theta$ and $\rho$ are localized around this ring, the boundary of the warp bubble.
@@
Ordinary gravity in three dimensions is special, it is dynamically trivial. Effects of localized sources are on the global properties whereas the outside metric is flat. For a point source of mass $M$ placed at the origin $r=0$, the spatial metric is a cone and the full metric is given by
@@
such that $v_{\infty}=\alpha^{2}\sqrt{\left(1-\frac{1}{\epsilon^{2}}\right)\,}$. Keeping in mind that the charts are invalid for $\left(\epsilon^{2}-\alpha^{2}\right)<0$, there are two natural choices for the parameter $\epsilon$ to realize an embedding of the warp drive in these defect backgrounds.

### v2

@@ Abstract
We extend the construction of Alcubierre–Natário class of warp drives to an infinite class of spacetimes with similar properties. This is achieved by utilising the Martel–Poisson charts which closely resembles the Weak Painlevé–Gullstrand form for various background metrics (Mink, AdS, dS). The highlight of this construction is the non–flat intrinsic metric which in three dimensional spacetimes introduce conical singularities at the origin and in higher dimensions generates non–zero Ricci scalar for the spatial hypersurfaces away from the origin. We analyse the expansion/contraction of space and the (NEC) violations associated with these warp drives and find interesting scalings due to the global imprints of the conical defects. Other properties like tilting of light cones, event horizons and several generalisations are also discussed.
@@
Warp Drive Conical defects Martel–Poisson ChartsPainlevé–Gullstrand NEC
@@ 1 Introduction
In this paper we have generalised the construction of Alcubierre–Natário like warp drive in three dimensional Minkowski spacetime by leveraging the Martel–Poisson (MP) family of charts Martel_2001 of which the Painlevé–Gullstrand (PG) coordinates are a special case. This construction has been extended to $\text{AdS}_{3}$ and dS backgrounds along with their conical defect counterparts. A generic key feature is the appearance of non–flat intrinsic metric with conical singularities which affect the global properties of the spacetimes. Another important realisation comes from the flow velocities of the background metrics where the expansion and $(-ve)$ energy density of the warp drive is considerably reduced if the drive travels along the background flow velocities. This will be useful in analog gravity setups which imitate the warp metric Fischer_2003 ; finazzi2012analogue and can be studied experimentally PhysRevLett.105.240401 . Throughout the paper we have used natural units $G=c=1$ and metric signature (,+,+,)(-,+,+$\cdots$)( - , + , + , ⋯ ).
The rest of the paper is organised as follows. In section 2 we review and extend the construction of MP–charts for spacetimes of the form (1) in $D$ dimensions. Interpretation of the flow coordinates are discussed including the special case of PG–coordinates. Section 3 highlights the conical singularities of spatial hypersurfaces in three dimensional Minkowski and AdS spacetimes and their conical defect cousins. Regularisation of the singularities and embedding of the warp drive in these spacetimes are discussed. Various properties, like tilting of light cones, event horizons and the violation of (NEC) along with total ($-ve$) energy estimates are explored. We conclude in section 4 with comments on possible explorations to higher dimensions, analogue gravity, quantum effects and optimisations. The paper ends with the relevant references.
@@ 2 Martel–Poisson charts
where $d\Omega_{(D-2)}^{2}$ is the line element on an unit $(D-2)$ sphere and $f(r)=(1+\tilde{f}(r))$  [^1] See later sections 3 and 4 for generalisations.. The main idea is to describe the spacetime in terms of radially infalling or outgoing timelike geodesic observers with D–velocities $u^{\mu}=\frac{dx^{\mu}}{d\tau}=(\dot{t},\dot{r},0,\cdots,0)$ where $\tau$ is the proper time along the geodesic. For static spacetimes, the Killing vector $\xi=(\partial/\partial t)$ leads to a conserved energy per unit mass $\epsilon=-\xi_{\mu}u^{\mu}=f\dot{t}$ and utilising the normalisation of the D–velocities $u_{\mu}u^{\mu}=-f\dot{t}^{2}+(\dot{r}^{2}/f)=-1$, we can finally write
@@
where the ($-ve$) and ($+ve$) signs are for the infalling and outgoing geodesics respectively.
So far, we have not established that these D–velocities are indeed geodesics. We demand that this congruence of D–velocities are hypersurface orthogonal to spacelike surfaces $T(t,r)$ = constant i.e. $u_{\mu}=-(\partial T/\partial x^{\mu})$ where $T$ is now the proper time for these free falling (geodesic) observers. Redefining the time coordinate as  [^2] Note that $T$ is an exact differential i.e. $\frac{\partial\epsilon}{\partial r}=\frac{\partial\left(\frac{\sqrt{\epsilon^{2}-f}}{f}\right)}{\partial t}\,$.
@@
Note that except for the case $\epsilon=1$, the induced metric on the hypersurface is not Ricci flat. In fact, as we shall see later the scalar curvature and other curvature invariants of the induced metric blows up at $r=0$.
@@
The observers can start at infinity but now with some initial velocity $v_{\infty}$ w.r.t. the time coordinate $t$ such that $\epsilon=1/\sqrt{1-v_{\infty}^{2}}\geq 1$ Again, zero initial velocity gives the PG–coordinates for the Schwarzschild black holes. There is another interesting limit where we can set $v_{\infty}=1$ i.e. the speed of light. Scaling $\tilde{T}=(T/\epsilon)$ in (3) and then taking the limit $\epsilon\rightarrow\infty$ gives the infalling Eddington–Finkelstein (EF) coordinates Faraoni2020WhenPC . Of course, it is also possible to have finite radius non–zero initial velocity observers but for the remainder of this paper we shall focus on the initial velocity interpretation for the parameter $\epsilon$.
@@ 3 Warp Drive in various spacetimes
While the original Alcubierre warp drive is a deformation of the Minkowski spacetime, we shall now discuss embedding the drive in various spherically symmetric static spacetimes, see also garattini2023black ; garattini2024wormholewarp . Our focus will be on the minimal non–trivial dimension $D=3$ from which it is relatively clear how to generalise for higher dimensions. Three dimensions is also special as it admits Conical defect solutions which change the global properties of the spacetime while retaining the local ones.
@@ 3.1 Minkowski spacetime
The MP–charts for the Minkowski spacetime in three dimensions with a tuneable parameter $\epsilon$ takes the form,
@@
where the radially infalling geodesic observers have the velocity $v_{\infty}=\sqrt{1-1/\epsilon^{2}}$ at spatial infinity. Though this metric has non–vanishing Christoffel symbols, it is just an artefact of the coordinate choice and the curvature tensors vanish.
@@
It is instructive to analyse the induced metric on the $T$ = constant hypersurfaces,
@@
We can however, introduce a regularisation by rolling off the cone tip. The simplest one corresponds to changing $\mathcal{C}_{\epsilon}$ to a hyperbolic space with a tuneable regularisation parameter $a$,
@@
To evaluate the curvature scalar, we first evaluate the integral curvature on the regularised hyperbolic space in the $a\rightarrow 0$ limit,
@@
where the second integral is zero. In a more general setting one can embed the conical defect in a non Ricci flat two dimensional manifold and the second integral would collect the curvature around the defect. We note that since only the singular point $\tilde{r}=0$ can give rise to the first term in (12), it is best to introduce a local representation for the curvature on $C_{\epsilon}\,$,
@@
We can embed the Alcubierre warp drive in the infalling Minkowski spacetime background (7). In Bobrick_2021 , various warp drives discussed in the literature are shown to be related to each other by coordinate transformations. It is crucial to note that the embedding discussed in this paper is not a coordinate transformations of the warp drive in the usual $\epsilon=1$ charts. Therefore, it represents an infinite class of different embeddings parameterised by $\epsilon$ or equivalently $v_{\infty}$.
@@
where $N$ is the lapse, $N^{i}$ are the shift vectors and $\gamma_{ij}$ are the components of the induced metric on the $T$ = constant spacelike hypersurfaces. For the metric as in (7), the embedding is as follows,
@@
where for three dimensions, $r=\sqrt{x^{2}+y^{2}}$ is the radial coordinate, $r_{s}=\sqrt{(x-x_{0}(T))^{2}+y^{2}}$ is the distance from the centre of the drive which is moving along the ($+ve$) $x$–direction with constant velocity $v$. Here, the form function $f(r_{s})$ decides the size and shape of the warp drive bubble, its exact form is of little consequence (unless one is optimising) as long as it has the value $f(0)=1$ at the centre of the bubble, $f\rightarrow 0$ as $r\rightarrow\infty$ outside the bubble and drops sharply from $1$ to 00 at the wall of the bubble at some distance $R$ from the centre of the bubble  [^3] One such specific function popular in the literature is $f(r_{s})=\frac{\tanh[\sigma(r_{s}+R)]-\tanh[\sigma(r_{s}-R)]}{2\tanh(\sigma R)}\,,$ (19) where $\sigma$ has the interpretation of being inversely proportional to the bubble wall thickness and can be taken to be large.. The choice of shift vectors $N^{i}$ are inspired by ellis2004warp , where the flow velocity is that of the warp drive inside the bubble and that of the background spacetime outside the bubble. The Eulerian observers with geodesics $u_{\mu}=(-1,0,0)$ are still in free fall and observe no time dilation. Hence, the above metric can be considered as a warp drive different from the Natário type Nat_rio_2002 . We assume the warp drive to be a test particle moving along the timelike curve $x=x_{0}(T)=vT$, regardless of the value of $v(T)$. One can verify that the proper time along this curve equals the coordinate time. Note that here, the induced metric is not flat and as discussed earlier will exhibit conical singularity at the origin. Warp drives with non flat $\gamma_{ij}$ have been discussed earlier in Broeck_1999 ; garattini2024wormholewarp . As expected, for $\epsilon=1$, we get back the original Alcubierre warp drive Alcubierre_1994 .
For the metric (7) which is just the $\epsilon\geq 1$, MP–chart for the Minkowski spacetime, we see that the light cones in the $r$–direction are $dr/dT=\pm 1-v_{\infty}$ i.e. they have a tilt proportional to $v_{\infty}$. This carries over to the generic form of the warp drives (14), only now the light cone along the $x$–direction has the tilt which can be a complicated function of the spatial metric and the lapse vectors, but at large $r$ takes a simplified form,
$\frac{dx}{dT}=\pm 1-(1-f(r_{s}))v_{\infty}+v\sqrt{(1-v^{2}_{\infty})}f(r_{s})\,.$ (20)
The warp drives travels along a timelike path inside this locally tilted light cones without violating anything in relativity, as the bubble is just a specific choice of the curvature of spacetime itself and not a massive object obeying the geodesic equation. Globally, the path taken by the warp drives can be spacelike for high enough warp velocity $v$.
We can also analyse the event horizons for these warp drives in a similar fashion. In order to modify the geometry of the warp bubble from inside i.e. to accelerate or to stop the spaceship, the crew must be able to causally influence the bubble wall in front of them by sending say a photon forward towards the bubble wall. In the reference frame of an observer on the spaceship, $x^{\prime}\equiv x-x_{0}(T)$, the forward photon has the trajectory,
$\frac{dx^{\prime}}{dT}=1-(1-f(r_{s}))\left(v_{\infty}+v\sqrt{(1-v^{2}_{\infty})}\right)\,.$ (21)
At the centre of the bubble $f(r_{s})=1$ and photon travels with the speed of light but if $\left(v_{\infty}+v\sqrt{(1-v^{2}_{\infty})}\right)>1$, then at some point inside the bubble $x=x_{h}$ where $f(z_{h})=1-\frac{1}{\left(v_{\infty}+v\sqrt{(1-v^{2}_{\infty})}\right)}$, the photon stops and remain forever at that point. Thus, there is a horizon at $x=x_{h}$ such that the outer edge of the bubble is outside the causal future $\mathcal{J}^{+}$ of the spaceship. A warp drive beyond a certain velocity can’t be controlled, though in principle it is possible that the entire trajectory of the bubble may be constructed beforehand. It is relatively straightforward to extend these arguments to the conical defect backgrounds and the warp drives in AdS/dS spacetimes discussed in the later sections.
@@
Few comments on the distance $r_{s}$ and the induced metric $\gamma_{ij}$ are in order. We can always choose $r_{s}$ to be the coordinate distance $r_{s}=\sqrt{(x-x_{0}(T))^{2}+y^{2}}$ from the center of the warp drive. In general, the induced metric $\gamma_{ij}$ is not $\delta_{ij}$, therefore the size & shape of the bubble inside the boundary $r_{s}=R$ will evolve over time. Even without the warp drive, the metric in (7) has non–zero expansion/contraction and shear for a spacelike patch orthogonal to the congruence of Eulerian geodesic observers  [^4] Non zero rotation is not possible as Eulerian geodesics are hypersurface orthogonal.. Our focus will be on the large $x_{0}$ and $(\epsilon-1)$ small limit where these background expansion and shear are at check. However, there are other choices, either $r_{s}=\sqrt{\gamma_{ij}(x-x_{0}(T))^{i}(x-x_{0}(T))^{j}}$ or $r_{s}=\text{Geodesic}(x,x_{0}(T))$ where $\text{Geodesic}(x,x_{0}(T))$ is the geodesic distance between the center of the warp drive $(x_{0}(T),0)$ and some point $(x,y)$. Another possible choice for the induced metric is $\gamma_{ij}=\delta_{ij}+(1-f(r_{s}))\left(\frac{1}{\epsilon^{2}}-1\right)\frac{x^{i}x^{j}}{r^{2}}\,$, an interpolating metric which is $\delta_{ij}$ inside the bubble and $\gamma_{ij}$ as in (17) outside  [^5] It is difficult to implement the other choices of $r_{s}$ with this choice of induced metric as now the definition of $r_{s}$ depends on $r_{s}$ itself through the function $f(r_{s})$. This puts severe constraints on possible $f(r_{s})$, which we checked to be incompatible with the desired properties of $f(r_{s})$.. We will discuss some of these cases in A.
@@
With our first choice of $r_{s}$ and the induced metric (17) we shall now discuss the movement of the warp bubble due to the expansion and contraction of the space around it. The space expands behind the warp bubble and contracts in front of it, thereby pushing the bubble forward (and maybe a spaceship inside it) at arbitrary velocity $v$  [^6] The same mechanism also applies to the the expanding universe, wherein galaxies move faster than light w.r.t. each other due to the expansion of the space itself.. We shall also discuss the energy density as measured by the Eulerian observers i.e. $\rho=T_{\mu\nu}u^{\mu}u^{\nu}$ where $T^{\mu\nu}$ is the stress–energy tensor of the metric and $u^{\mu}=(1,-N^{x},-N^{y})$ are the 3–velocities.
@@
where $D_{i}$ are the covariant derivatives w.r.t. the induced metric $\gamma_{ij}$. The expansion is the trace $\theta=K^{i}_{i}$. The Einstein tensor $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}$ along with the Gauss–Codazzi relations Gourgoulhon:2012ffd which connects curvature tensors of the metric to curvatures on the hypersurface(s) give,
@@
$D^{k}K_{ik}-D_{i}K$ $=8\pi f_{i}$
where $\rho$ is the matter energy density and $f_{i}$ are the matter momentum density and $\Lambda$ is the cosmological constant.
For the background metric in (7) without the warp drive, the extrinsic curvature and expansion away from the conical singularity at $r=0$ are given by  [^7] The regularisation mentioned in (11) needs modification as now we want ${}^{(2)}R+\left(K^{2}-K^{ij}K_{ij}\right)$ to vanish at $r=0$ as $\rho=\Lambda=0$. The regularised metric around $r=0$ with $0\leq\phi<2\pi$ is given by $ds^{2}=-dT^{2}+\frac{1}{\epsilon^{2}}\left(A(r)\,dr+\sqrt{\epsilon^{2}-1}\,B(r)\,dT\right)^{2}+r^{2}d\phi^{2}$ $A(r)=\sqrt{\frac{a^{2}\epsilon^{4}+r^{2}}{a^{2}\epsilon^{2}+r^{2}}}\,\,\,\,\,B(r)=\frac{\sqrt{\epsilon^{2}(1-2\ln A(r))-1}}{\sqrt{\epsilon^{2}-1}}$ (24)
@@
Closed form expressions for the $K_{ij}$ and $\rho$ can be obtained for the metric with the warp drive centred at $x_{0}$. However, we are interested in the limit $x_{0}\gg(x-x_{0})\approx y\approx r_{s}\,$, where the warp drive is sufficiently far way from the centre $r=0$. We shall also take $v_{\infty}\ll 1$ for the Eulerian observers. To $\mathcal{O}\left(\frac{1}{r}\right)$ and $\mathcal{O}(v_{\infty})$, the expressions are as follows  [^8] Here, ${}^{(2)}R=0$ on the hypersurfaces, except at $r=0$. This is not true in $D\geq 4$ dimensions, thereby opening up new possibilities to minimise the energy density.
@@
At first, we note that for $v_{\infty}=0$, the expressions falls back to the usual Alcubierre warp drive in the Minkowski spacetime. As expected, the $1/r$ zeroth order expressions for both $\theta$ and $\rho$ have the combination $(v+v_{\infty})$ for velocity. In these river flow coordinates, the background metric is flowing like a fluid along the radially inward direction. However, the warp drive is flowing against this “fluid” in the $(+ve)$ $x$–direction, i.e. it is flowing with velocity $(v+v_{\infty})$ w.r.t. the fluid. Therefore, it is best to move along the flow to minimise the peak expansions/contractions and $(-ve)$ energy density. In fact, for a free falling warp drive to zeroth order they are zero but picks up a ($-ve$) contribution at $\mathcal{O}\left(\frac{1}{r^{2}}\right)$. We also note that, the first $\mathcal{O}\left(\frac{1}{r}\right)$ term in $\theta$ is that of the background metric and maybe subtracted, the other one is proportional to $(y/r)$ and hence very small. As $f^{\prime}(r)<0$, behind the bubble $\theta>0$ and the space is expanding, while at the front $\theta<0$ and the space is contracting. By construction $f^{\prime}(r_{s})$ is non zero only along a very small width of the ring $r_{s}=R$. Hence, the effects of the warp drive in both $\theta$ and $\rho$ are localised around this ring, the boundary of the warp bubble.
@@
Ordinary gravity in three dimensions is special, it is dynamically trivial. Effects of localised sources are on the global properties whereas the outside metric is flat. For a point source of mass $M$ placed at the origin $r=0$, the spatial metric is a cone and the full metric is given by
@@
such that $v_{\infty}=\alpha^{2}\sqrt{\left(1-\frac{1}{\epsilon^{2}}\right)\,}$. Keeping in mind that the charts are invalid for $\left(\epsilon^{2}-\alpha^{2}\right)<0$, there are two natural choices for the parameter $\epsilon$ to realise an embedding of the warp drive in these defect backgrounds.